OpenStudy (anonymous):

How many solutions does this equation have 2(2x+5)=4(x+3)

OpenStudy (texaschic101):

2(2x + 5) = 4(x + 3) ---- distribute through the parenthesis 4x + 10 = 4x + 3 --- subtract 4x from both sides 4x - 4x + 10 = 4x - 4x + 3 -- combine like terms 10 = 3 (lol....nope, not true) Therefore, there are 0 solutions to this problem